40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Thu Oct 21 15:32:32 2021, Aidan, Radhika, Laser, 2um Photodiodes, Improved measurement of QE on photodiodes ~89% at 140K IMG_5305.jpg
    Reply  Thu Oct 21 16:52:26 2021, rana, Laser, 2um Photodiodes, Improved measurement of QE on photodiodes ~89% at 140K 
       Reply  Thu Oct 21 17:48:33 2021, Aidan, Laser, 2um Photodiodes, Improved measurement of QE on photodiodes ~89% at 140K 
          Reply  Fri Oct 22 09:20:13 2021, Aidan, Laser, 2um Photodiodes, Improved measurement of QE on photodiodes ~89% at 140K 
          Reply  Fri Oct 22 12:04:13 2021, rana, Laser, 2um Photodiodes, Improved measurement of QE on photodiodes ~89% at 140K 
             Reply  Fri Oct 22 13:31:36 2021, Chris, Laser, 2um Photodiodes, Improved measurement of QE on photodiodes ~89% at 140K 
Message ID: 2680     Entry time: Thu Oct 21 15:32:32 2021     Reply to this: 2681
Author: Aidan, Radhika 
Type: Laser 
Category: 2um Photodiodes 
Subject: Improved measurement of QE on photodiodes ~89% at 140K 

[Aidan, Radhika]

We turned off the heater and the cryocooler this morning to around 11:30AM when the temperature of the diode was around 123K and are letting it gradually warm up.

Through the next 30K, we experimented with different distances between the fiber output and the collimating lens. Bias voltage always set to 1000mV. Laser diode current was set manually on the controller (the input from the DAC was unplugged as this is a little noisy). 

  1. We increased the distance between collimator and the lens by 1.4mm (from stage reading of 9.41mm to 10.8mm) and there was a small increase in the response (PD output/REF PD output)
  2. For each new setting, we set the laser diode current to 100mA and run the maximize script on the piezo mirror, adjusting the alignment to maximize the output power.
  3. We then stepped up the laser diode current in steps of 10mA from 25mA to 95mA and one last measurement at 101mA. The PD response dropped by 30-40% through this range.

It looked like the optimum translation stage setting was 9.75mm - however, i discovered something very interesting ...

If you run the maximize power script at 100mA to the laser diode, then drop the laser current to 30mA and rerun the script, you find that there is a different optimum alignment. This means that the output beam shape/pointing is power dependent. In other words, the output of the fiber is not properly mode-cleaned by the 2m patch cord we have.

Switching to 25mA, I optimized the alignment and continued exploring the optimum translation stage position. Dropping the stage position to 8.0mm maximized the response (at 25mA). Note that the code maximizes an EPICS channel called C4:TST-PD_RESPONSE which is (JPL_PD - DarkV)/REFPD. The reference PD filter bank has an offset applied so OUT16 has a mean value of zero when the laser is off. DarkV for the JPL PD is the PD voltage when the laser is off and this was manually updated every 5-10 minutes or so. C4:TST_PD_RESPONSE is not stored but the JPL_PD and REFPD channels are stored in frames.

On 27-Sept, I measured the ratio of power incident on the JPL PD to the voltage output from the REF PD: dP/dV = 9.3E-4 W/V. The JPL PD DC photocurrent sees a transimpedance gain of 1000Ohm. Therefore, QE is calculated using the following formula:

QE = (RESPONSE)*(1E-3/9.3E-4)*  h*c/(e*lambda) = RESPONSE*0.667

Using this calculation and a peak response value of 1.334 at about 145K, the peak QE was estimated to be about 89%. An error analysis is needed on this. And we need to figure out how to get a better output beam shape from the optical fiber (use a really long fiber?)

Note that the translation stage reading of 8.0mm corresponds to a fiber holder to lens mount distance of 30.9mm

 

Attachment 1: IMG_5305.jpg  2.106 MB  Uploaded Thu Oct 21 17:07:35 2021  | Hide | Hide all
IMG_5305.jpg
ELOG V3.1.3-